# DreamSim: Learning New Dimensions of Human Visual Similarity using Synthetic Data

### v1

@@ 6 Applications
Image retrieval.
We use our metric for image retrieval on two datasets: ImageNet-R hendrycks2021many , which contains 30K images of various renditions of ImageNet categories (e.g., sketches, cartoons), and a random subset of 30K images from MS-COCO lin2014microsoft , which contains diverse natural photos. Given a query image, we compute its similarity to the entire dataset and present the top-3 nearest images under each metric in Fig. 8. As can be seen, our metric finds meaningful neighbors, demonstrating that it can generalize to diverse image domains. Compared to the baselines, which are either dominated by low-level similarity (e.g., background color in LPIPS/DIST), or higher-level similarity (e.g., semantic class in OpenCLIP), our retrieved images exhibit similarity across a continuum of visual traits.
@@
Feature inversion.
To better understand the information captured by our metric, we apply existing feature inversion techniques, where the aim is to optimize for an image that is similar to a target image under some metric. We compare our metric to the DINO and OpenCLIP embeddings, as well as to an ensemble embedding that simply concatenates all features without finetuning.
Fig. 9 shows the inversion results under a range of image priors. Under vanilla optimization, i.e., not using any image prior mahendran2015understanding ; olah2017feature , our metric’s results exhibit higher fidelity to the original colors, shapes, and semantics. Using Deep Image Prior ulyanov2018deep ; tumanyan2022splicing , our metric reveals better global layout preservation and semantics. This is also evident when using our metric to guide recent generative models dhariwal2021diffusion , resulting in higher quality image samples that resemble the semantic concepts and appearance of the target image.
Consistent with Sec. 5.2 and the retrieval results, these visualizations demonstrate our metric’s sensitivity to color, layout, and semantics. The increased quality of our fine-tuned embeddings compared to just ensembling shows the effectiveness of tuning on our human-aligned dataset. See SM for more examples.
Figure 9: Feature inversion across different metrics and image priors. Given a target image, we optimize for an image, where the objective is to match the target image embedding of a given backbone. Without any image prior (Optimization), our metric recovers better the color, shape and semantics of the target image. With a weak image prior ulyanov2018deep ; tumanyan2022splicing (DIP Inversion), our metric is better able to reproduce scene structure and semantics. Using a diffusion model dhariwal2021diffusion as a strong prior, our metric better captures overall semantics and scene appearance.
@@ 7 Discussion
We expand the fundamental task of measuring perceptual image similarity to encompass factors beyond mere low-level similarity, spanning multiple notions of similarities (e.g., object poses, colors, shapes, and camera angles). By harnessing a state-of-the-art generative model, we overcome the lack of suitable data and introduce a new dataset consisting of human-judged synthetic image triplets. We evaluate candidate embeddings and propose a new similarity metric that better aligns with human judgments, demonstrating its generalization and effectiveness compared to existing metrics. We note that by using Stable Diffusion (SD) rombach2022high , our benchmark is exposed to potential preexisting biases and sensitive content in the model. As such, we generate our images with a pre-defined set of single-word categories, while avoiding human faces. Our perceptual model is also finetuned from existing pre-trained backbones, and thus may also inherit prior errors and biases. However, our work is a step in aligning representations to human preferences, and we believe our work can trigger further research, leveraging advances in generative image models for perceptual studies.
Acknowledgments We thank Jim DiCarlo, Liad Mudrik, Nitzan Censor for fruitful discussions throughout the project. Additionally, we thank Narek Tumanyan for his insightful comments over the course of the project. Finally, we thank Michelle Li for proofreading sections of this paper and offering helpful comments. This work was supported by a Packard Fellowship to P.I., by the Israeli Science Foundation (grant 2303/20) to T.D., Meta PhD Fellowship to L.C., and NSF GRFP Fellowship to S.S.
@@
In the supplemental document, Sec. A contains additional methodological details on dataset collection and model training, Sec. B provides more analyses of our model, and Sec. C discusses the broader impact of our work, limitations, and licensing. Additional qualitative results, such as additional triplets from our dataset and visualizations of image retrieval and image reconstruction, are available on our project page.
@@ A.1 AMT Details
Users are shown four images for 500 ms each with a 1 second gap inbetween, and are then prompted to answer the two questions in Figure 10 (right). They are given feedback and a score, though these have no bearing on the JND results themselves. Each user completes 2 practice tasks, 24 tasks with “different" pairs, and 12 tasks with “same" pairs, averaging 10 minutes for the entire assignment.
@@
Figure 10: User interface for AMT studies. (Left) One image triplet shown to a user in 2AFC, who is prompted to pick "Image A" or "Image B". (right) In each JND task, users are shown a sequence of images and asked whether the image pairs were identical. Upon answering, they are given the correct answers.
@@
Cost breakdown. Across 10 rounds of 2AFC studies, we show users 477,964 triplets (see Table 2 for the full breakdown). Each user is paid \$0.50 for one assignment consisting of 50 triplets, averaging \$10.00/hr. In total, we pay users \$4779.64 to collect 2AFC judgments on our dataset.
We run JND on image triplets in our test set that received >5 2AFC judgments (1,824 triplets total). Each user is paid \$2.00 for one assignment consisting of 48 image pairs, averaging \$12.00/hr. In total, we pay users \$156.00 to collect JND judgments on our test set.
@@ A.2 Additional Dataset Details
2AFC task. Following Sec. 3.2 in the main text, we filter the images over 10 rounds to obtain cognitively impenetrable triplets where humans tend to vote the same way despite various differences between the images. Statistics for each round of filtering is reported in Tab. 2, which leaves us with roughly 20% of the original triplets containing unanimous votes. We discard all votes from any turker who fails the sentinel task. As a result, not all triplets have the same number of votes. The resulting sizes of the train, validation, and test splits and additional statistics on each split are reported in Tab. 3-top. In practice, we further discard triplets with five or fewer unanimous votes.
JND task. The JND triplets are meant to capture the decision boundary where two different images are similar enough to be confused as identical, in the presence of a masking image. Each triplet is divided into two pairs – Ref vs. A and Ref vs. B. These pairs are presented to different turkers in different interleaving sequences, and we collect three judgments for each pair and take the majority vote among the three judgments, thus six judgments in total per triplet (Tab. 3-bottom).
@@
For all fine-tuned models (both Tuned - MLP and Tuned - LoRA) we use the NIGHTS training dataset, with an 80%-10%-10% split as described in Tab. 3, and images of size $768\times 768$ resized to $224\times 224$. We train on a single NVIDIA GeForce RTX 3090 or NVIDIA TITAN RTX GPU with an Adam optimizer, learning rate of $43e-$ 4, weight decay of 0, and batch size of 512 (non-ensemble models) and 16 (ensemble models). In MLP-tuned training, we use a width of 512. We tune the number of training epochs using the validation set; for the Tuned - LoRA ensemble model (DreamSim) we train for 6 epochs. For Tuned - LoRA models we use rank $r=16$, scaling $\alpha=0.5$, and dropout $p=0.3$. Training time is approximately 30 min/epoch for LoRA-tuned models, and 15 min/epoch for MLP-tuned models.
@@ B.1 Additional Evaluations
Full Evaluation on Large Vision Models. In Sec. 5.1 and Fig. 5 in the main text we report results using the best-performing setting of various large vision model backbones. Tab. 4 shows the experimental variation over multiple runs, in which the LoRA variation consistently outperforms the MLP variation. Tab. 5 evaluates additional model settings, spanning different ViT model sizes, patch sizes, and strides.
@@
As some models are adapted from backbones trained on ImageNet deng2009imagenet (including the prior learned metrics and DINO), we split our dataset into categories contained in ImageNet and those not in ImageNet, and evaluate alignment on each split. Performance on both splits is highly correlated (Fig. 12), suggesting that the notions of visual similarity are related regardless of whether or not the triplet was generated from an ImageNet category, and whether or not the model was trained only on ImageNet.
@@
Alignment with low-level features. In Sec. 5.2 and Fig. 7 of the main text we report results on the alignment between our metric, OpenCLIP, DINO, LPIPS, and DISTS with low-level and semantic metrics for the COCO dataset. In Tab. 6 we also report additional, fine-grained results for COCO triplets. We use CarveKit carvekit to segment out the foreground and background of each image, and then breakdown how well each metric agrees with RGB color histogram similarity, luminance histogram similarity, and depth map distance, for foreground, background, and the full image. For color histograms we use 32 bins for each channel, and for luminance histograms we use 10 bins.
@@
We also examine semantic features. For each image, we find the percentage of area that each semantic category occupies, and then compute the alignment between the difference in area for each category and perceptual metrics. Note that when the difference in area is the same for both pairs in a triplet, it is counted as 50% alignment. When the difference in area is smaller for the pair chosen by the perceptual metric, it is counted as 100% alignment (and 0% in the case of disagreement). In Tab. 6 we show the five semantic categories most aligned with our metric. Our metric has a 55% alignment score with the "people" category, however does not seem to align well above chance for other categories.
@@
In Tab. 7 we show alignment with low-level metrics for our dataset (which does not have semantic annotations). On our dataset there is higher alignment with color, luminance, and depth across all metrics, as compared to COCO triplets. This is likely because the images in each of our dataset’s triplets all share the same semantic category, making lower-level features more important than for the randomly-chosen COCO triplets. Our model aligns significantly better with foreground metrics – particularly foreground color – whereas LPIPS aligns slightly better with background.
@@
Dimensionality reduction with PCA. Our model consists of the concatenation of the DINO, OpenCLIP, and CLIP backbones, and therefore uses a higher-dimensional feature space to compute similarity compared to each of these models independently. To investigate whether the increased dimensionality is critical for improving human alignment, we ablate feature dimensions by applying PCA, taking a certain number of the top components, as seen in Fig. 14 and Tab. 8. We can achieve comparable performance using just 500 of the top components, similar to the 512 dimensions of the CLIP and OpenCLIP embedding outputs, suggesting that the improved alignment is not just due to the higher-dimensional feature space used to compute similarity, but rather the additional capacity and model priors obtained from ensembling different models.

### v3

@@ 6 Applications
Image retrieval. We use our metric for image retrieval on two datasets: ImageNet-R hendrycks2021many , which contains 30K images of various renditions of ImageNet categories (e.g., sketches, cartoons), and a random subset of 30K images from MS-COCO lin2014microsoft , which contains diverse natural photos. Given a query image, we compute its similarity to the entire dataset and present the top-3 nearest images under each metric in Figure 8. As can be seen, our metric finds meaningful neighbors, demonstrating that it can generalize to diverse image domains. We compare to the best performing ViT embedding methods (OpenCLIP, DINO) and prior learned metrics (LPIPS, DISTS), as evaluated in Section 5.1. Compared to the baselines, which are either dominated by low-level similarity (e.g., background color in LPIPS/DISTS), or higher-level similarity (e.g., semantic class in OpenCLIP), our retrieved images exhibit similarity across a continuum of visual traits. See Section B.1 for more results and analysis on sketch-to-photo and photo-to-sketch domains.
@@
Figure 9: User preferences for image retrieval results by metric. We conduct a user study that collects preferences for retrieval results from LPIPS, DISTS, DINO, OpenCLIP, and DreamSim. We visualize one standard deviation above and below each bar. On ImageNet-R (top, orange), our metric is preferred by users 36.8% of the time, followed by OpenCLIP (28.7%). Similarly, on COCO (bottom, blue), users prefer our metric 33.6% of the time, with DINO being the second choice (28.1%).
Feature inversion. To better understand the information captured by our metric, we apply existing feature inversion techniques, where the aim is to optimize for an image that is similar to a target image under some metric. We compare our metric to the DINO and OpenCLIP embeddings, as well as to an ensemble embedding that simply concatenates all features without finetuning.
Figure 10 shows the inversion results under a range of image priors. Under vanilla optimization, i.e., not using any image prior mahendran2015understanding ; olah2017feature , our metric’s results exhibit higher fidelity to the original colors, shapes, and semantics. Using Deep Image Prior ulyanov2018deep ; tumanyan2022splicing , our metric reveals better global layout preservation and semantics. This is also evident when using our metric to guide recent generative models dhariwal2021diffusion , resulting in higher quality image samples that resemble the semantic concepts and appearance of the target image.
Consistent with Section 5.2 and the retrieval results, these visualizations demonstrate our metric’s sensitivity to color, layout, and semantics. The increased quality of our fine-tuned embeddings compared to just ensembling shows the effectiveness of tuning on our human-aligned dataset. See SM for more examples, and Section A.4 for details on feature inversion methodology.
k-NN Classification. We evaluate our metric as a k-Nearest Neighbors classifier, which requires the retrieval of images that are both visually and semantically relevant to a given query. We find that our metric outperforms all baselines on the ObjectNet objectnet dataset and performs competitively with DINO on the ImageNet100 tian2020contrastive dataset. For full details and results, refer to SM Section B.1.
Figure 10: Feature inversion across different metrics and image priors. Given a target image, we optimize for an image, where the objective is to match the target image embedding of a given backbone. Without any image prior (Optimization), our metric better recovers the color, shape and semantics of the target image. With a weak image prior ulyanov2018deep ; tumanyan2022splicing (DIP Inversion), our metric is able to reproduce scene structure and semantics. Using a diffusion model dhariwal2021diffusion as a strong prior, our metric better captures overall semantics and scene appearance.
@@ 7 Discussion
We expand the fundamental task of measuring perceptual image similarity to encompass factors beyond mere low-level similarity, spanning multiple notions of similarities (e.g., object poses, colors, shapes, and camera angles). By harnessing a state-of-the-art generative model, we overcome the lack of suitable data and introduce a new dataset consisting of human-judged synthetic image triplets. We evaluate candidate embeddings and propose a new similarity metric that better aligns with human judgments, demonstrating its generalization and effectiveness compared to existing metrics. We note that by using Stable Diffusion (SD) rombach2022high , our benchmark is exposed to potential preexisting biases and sensitive content in the model. Our perceptual model is also finetuned from existing pre-trained backbones, and thus may also inherit prior errors and biases. However, our work is a step in aligning representations to human preferences, and we believe our work can inspire further research in leveraging advances in generative image models for perceptual studies.
Acknowledgments. We thank Jim DiCarlo, Liad Mudrik, Nitzan Censor, Michelle Li, and Narek Tumanyan for fruitful discussions throughout the project. This work was supported by a Packard Fellowship to P.I., Israeli Science Foundation (grant 2303/20) to T.D., Meta PhD Fellowship to L.C., and NSF GRFP Fellowship to S.S.
@@
In the supplemental materials, Sec. A contains additional methodological details on dataset collection and model training, Sec. B provides more analyses of our model, and Sec. C discusses the broader impact of our work, limitations, and licensing. Additional qualitative results, such as additional triplets from our dataset and visualizations of image retrieval and image reconstruction, are available on our project page.
@@ A.1 AMT Details
Users are shown four images for 500 ms each with a 1 second gap inbetween, and are then prompted to answer the two questions in Figure 11 (right). They are given feedback and a score, though these have no bearing on the JND results themselves. Each user completes 2 practice tasks, 24 tasks with “different” pairs, and 12 tasks with “same” pairs, averaging 10 minutes for the entire assignment.
@@
Figure 11: User interface for AMT studies. (Left) One image triplet shown to a user in 2AFC, who is prompted to pick “Image A” or “Image B”. (Right) In each JND task, users are shown a sequence of images and asked whether the image pairs were identical. Upon answering, they are given the correct answers.
@@
Cost breakdown. Across 10 rounds of 2AFC studies, we show users 477,964 triplet instances (see Table 2 for the full breakdown). Each user is paid \$0.50 for one assignment consisting of 50 triplets, averaging \$10.00/hr. In total, we pay users \$4779.64 to collect 2AFC judgments on our dataset.
We run JND on image triplets in our test set that received $\geq$6 2AFC judgments. Each user is paid \$2.00 for one assignment consisting of 48 image pairs, averaging \$12.00/hr. In total, we pay users \$156.00 to collect JND judgments on our test set.
@@ A.2 Additional Dataset Details
Filtering text prompts. As described in Sec. 3.1 we sample images using image labels drawn from popular image classification datasets. We perform basic filtering to avoid human categories that typically result in malformed generated images, such as “man” and “baby”. Thus, human faces only feature in a small percentage of our dataset (<0.5%), typically in categories such as “lab coat” and “bonnet” that are still associated with humans. We note that even with some human categories excluded, our model is still sensitive to the presence of people, as shown in Fig. 7 and Fig. 8. For a full list of categories used, refer to our GitHub page.
2AFC task. Following Sec. 3.2 in the main text, we filter the images over 10 rounds to obtain cognitively impenetrable triplets where humans tend to vote the same way despite various differences between the images. Statistics for each round of filtering is reported in Table 2, which leaves us with roughly 20% of the original triplets containing unanimous votes. We discard all votes from any worker who fails the sentinel task. As a result, not all triplets have the same number of votes. The resulting sizes of the train, validation, and test splits and additional statistics on each split are reported in Table 3 (top). In practice, we further discard triplets with five or fewer unanimous votes.
JND task. The JND triplets are meant to capture the decision boundary where two different images are similar enough to be confused as identical, in the presence of a masking image. Each triplet is divided into two pairs: Ref vs. A and Ref vs. B. These pairs are presented to different workers in different interleaving sequences. We collect three judgments for each pair and take the majority vote among the three judgments, thus collecting six judgments in total per triplet (Table 3 (bottom)).
@@
For all fine-tuned models (both Tuned - MLP and Tuned - LoRA) we use the NIGHTS training dataset, with an 80%-10%-10% split as described in Table 3, and images of size $768\times 768$ resized to $224\times 224$. We train on a single NVIDIA GeForce RTX 3090 or NVIDIA TITAN RTX GPU with an Adam optimizer, learning rate of $43e-$ 4, weight decay of 0, and batch size of 512 (non-ensemble models) and 16 (ensemble models). In MLP-tuned training, we use a width of 512. We tune the number of training epochs using the validation set; for the Tuned - LoRA ensemble model (DreamSim) we train for 6 epochs. For Tuned - LoRA models we use rank $r=16$, scaling $\alpha=0.5$, and dropout $p=0.3$. Training time is approximately 30 min/epoch for LoRA-tuned models, and 15 min/epoch for MLP-tuned models.
A.4 Feature Inversion
In Fig. 10 we show examples of using our metric to optimize for an image generation that is similar to some target image. We do so in three experiments, with increasingly strong generative priors.
Optimization. In this setup, we initialize an image randomly and update it iteratively using gradient descent with respect to the distance between it and the target image. Given a random crop of our working image and the target image we compute the distance between the two using a specific metric (DINO/OpenCLIP/Ensemble/DreamSim) and their corresponding embeddings. Note that in this setup there is no generator neural network.
Deep Image Prior. This technique involves beginning with a frozen random noise which serves as an input to a trainable U-Net. The U-Net is optimized iteratively by comparing the output of the U-Net with the target image using different metrics and their embeddings. This introduces a CNN prior on the generated image as described in the Deep Image Prior paper ulyanov2018deep .
Guided Diffusion. This technique involves using classifier guidance to guide the generation of a 256x256 unconditional ImageNet-trained diffusion model NEURIPS2021_49ad23d1 using an input prompt image. The guidance is done by computing the gradient of some loss function, given the estimated clean image and the target image, with respect to the current iteration. At every training step we compute the spherical distance loss between our model’s embeddings of 64 random crops of the input image and the synthesized image. We combine this distance with a total variation loss, following mullisClipGuided ; kcrowsonKDiffusion . The gradient of this loss is then added to the mean of the current iteration, which pushes the generated image to be closer to the target image.
@@ B.1 Additional Evaluations
Full Evaluation on Large Vision Models. In Section 5.1 and Figure 5 in the main text we report results using the best-performing setting of various large vision model backbones. Table 4 shows the experimental variation over multiple runs, in which the LoRA variation consistently outperforms the MLP variation. Table 5 evaluates additional model settings, spanning different ViT model sizes, patch sizes, and strides.
@@
As some models are adapted from backbones trained on ImageNet deng2009imagenet (including the prior learned metrics and DINO), we split our dataset into categories contained in ImageNet and those not in ImageNet, and evaluate alignment on each split. Performance on both splits is highly correlated (Figure 13), suggesting that the notions of visual similarity are related regardless of whether or not the triplet was generated from an ImageNet category, and whether or not the model was trained only on ImageNet.
@@
Alignment with low-level features. In Section 5.2 and Figure 7 of the main text we report results on the alignment between our metric, OpenCLIP, DINO, LPIPS, and DISTS with low-level and semantic metrics for the COCO dataset. In Table 6 we also report additional, fine-grained results for COCO triplets. We use CarveKit carvekit to segment out the foreground and background of each image, and then break down how well each metric agrees with RGB color histogram similarity, luminance histogram similarity, and depth map distance for foreground, background, and the full image. For color histograms we use 32 bins for each channel, and for luminance histograms we use 10 bins.
@@
We also examine semantic features. For each image, we find the percentage of area that each semantic category occupies, and then compute the alignment between the difference in area for each category and perceptual metrics. Note that when the difference in area is the same for both pairs in a triplet, it is counted as 50% alignment. When the difference in area is smaller for the pair chosen by the perceptual metric, it is counted as 100% alignment (and 0% in the case of disagreement). In Table 6 we show the five semantic categories most aligned with our metric. Our metric has a 55% alignment score with the “people” category, though does not align well above chance for other categories.
@@
In Table 7 we show alignment with low-level metrics for our dataset (which does not have semantic annotations). On our dataset there is higher alignment with color, luminance, and depth across all metrics, as compared to COCO triplets. This is likely because the images in each of our dataset’s triplets all share the same semantic category, making lower-level features more important than for the randomly-chosen COCO triplets. Our model aligns significantly better with foreground metrics – particularly foreground color – whereas LPIPS aligns slightly better with background.
Sketch-photo image retrieval. To further analyze perceptual metric performance on out-of-domain data, we evaluate sketch-to-photo and photo-to-sketch image retrieval on the Sketchy database sketchy2016 . First, we assign sketches as queries and photos as the search space. As seen in Figure 16 (left), DINO, DISTS, and LPIPS perform poorly on this task compared to OpenCLIP and DreamSim because they focus on color similarity, retrieving photos with predominantly white/gray backgrounds. This sometimes appears in DreamSim’s results as well, hindering performance (Figure 15, top-left). Without this failure case, we return impressive results (Fig. 15, top-right) focusing on the subject appearance. We also evaluate retrieval in the other direction, using photos as queries and sketches as the search space. As shown in Figure 15 (bottom), DreamSim outperforms all other metrics in retrieving the main object while remaining sensitive to its size and location in the image.
Figure 15: Sketch-photo nearest-neighbor retrieval. Image retrieval on the Sketchy database sketchy2016 , with a sketch as query (top) and photo as query (bottom). OpenCLIP and DreamSim best preserve image content, returning photos of the subject itself (e.g. church, top right) rather than focusing on the black-and-white sketch style in the query. Our metric also best localizes the subject, returning a small boat in the corner while OpenCLIP and DINO both miss this spatial detail (bottom right).
Figure 16: Sketch-photo retrieval recall scores. When using real photos as the queries, our model outperforms all methods across # nearest neighbors, obtaining a Recall@10 score of 0.32, compared to the next best model, OpenCLIP, with Recall@10 = 0.20. When using sketches as the queries, we find that our model is competitive with OpenCLIP and outperforms the remaining models.
Image Quality Assessment (IQA) benchmarks. We evaluate various metrics on the TID2013 tid2013 and KADID-10k kadid10k image quality assessment (IQA) benchmarks and report results in Table 8. Following standard IQA literature, we measure the Spearman Rank Order Correlation Coefficient (SROCC) between metric similarities and Mean Opinion Score (MOS). While LPIPS and FSIM remain best-performing overall, DreamSim is still competitive in IQA, outperforming most low-level metrics and base ViT models. Notably, our metric improves SROCC by $\sim 0.05$ after training on NIGHTS despite not having seen these low-level distortions. These results are consistent with those in Figure 14, where tuning on NIGHTS consistently improves human alignment on BAPPS.
TID2013 KADID-10k
Low-level PSNR* 0.687 0.676
SSIM 0.720 0.724
MS-SSIM 0.798 0.802
FSIM* 0.851 0.854
Prior learned LPIPS 0.838 0.883
DISTS 0.794 0.867
Base models DINO 0.722 0.753
OpenCLIP 0.764 0.841
Ensemble 0.757 0.812
Ours DreamSim 0.814 0.868
Table 8: Evaluation on IQA datasets. Alignment of various models on TID2013 tid2013 and KADID-10k kadid10k While LPIPS and FSIM are best-performing overall, DreamSim is still competitive in IQA despite not having been tuned on low-level distortions. $*{}^{*}$hile in some IQA literature PSNR/FSIM refer to computation on the luminance channel and PSNRc/FSIMc on color channels, we keep the metric names as-is and compute on color channels to be consistent with our PSNR labels elsewhere in the paper.
k-Nearest Neighbors Classification. We evaluate our model as a k-Nearest Neighbor (k-NN) classifier on the ObjectNet objectnet and ImageNet100 cmcTian datasets, both standard classification benchmarks. Results are shown in Figure 17. Similarly to our other applications, we also evaluate OpenCLIP, DINO, and LPIPS as baselines (we were unable to evaluate DISTS due to computational constraints).
For ImageNet100 (a 100-category subset of ImageNet), we evaluate on the validation set and use the training set as the candidate search space. We note that DINO was trained using self-supervision on ImageNet, and has been previously shown to achieve strong k-NN results on ImageNet. DINO performs best, however our model still achieves competitive performance. For ObjectNet, we use an 80:20 split to randomly partition the dataset into a search-space and test set. Our model outperforms both OpenCLIP and DINO on top-1 classification across different values of $k\in\{1,3,5,11,21,101,201\}$.
A common failure case for all three models is when the retrieved neighbors are visually similar to the query image but of a different category. We also note that ObjectNet is designed to differ from the ImageNet distribution, indicating that DreamSim may generalize well across different image distributions.
Figure 17: k-NN classification results. When evaluated as a k-NN classifier on the ObjectNet dataset, our model outperforms all methods across different values of $k$, with a maximum accuracy of 30.92%; the next best method, OpenCLIP, achieves 22.48%. When evaluating on ImageNet100, our model is competitive (1-2% difference) with DINO and outperforms other baselines.
@@
Dimensionality reduction with PCA. Our model consists of the concatenation of the DINO, OpenCLIP, and CLIP backbones, and therefore uses a higher-dimensional feature space to compute similarity compared to each of these models independently. To investigate whether the increased dimensionality is critical for improving human alignment, we ablate feature dimensions by applying PCA, taking a certain number of the top components, as seen in Figure 18 and Table 9. We can achieve comparable performance using just 500 of the top components, similar to the 512 dimensions of the CLIP and OpenCLIP embedding outputs, suggesting that the improved alignment is not just due to the higher-dimensional feature space used to compute similarity, but rather the additional capacity and model priors obtained from ensembling different models.